# Equal contribution    Project lead    Corresponding author

### v2

@@ 1 Introduction
At the core of TSPMs is their ability to recognize and generalize time series patterns inherent in time series data, enabling the model to uncover meaningful temporal structures and adapt to varying time series task scenarios. A line of research (Lai et al., 2018b; Zhao et al., 2017) has utilized recurrent neural networks (RNNs) to capture sequential patterns. However, these methods often struggle to capture long-term dependencies due to limitations like Markovian assumptions and inefficiencies. Temporal convolutional networks (TCNs) (Wu et al., 2023; Wang et al., 2023a; Liu et al., 2022a; Wang et al., 2024a) efficiently capture local patterns but face challenges with long-range dependencies (e.g., seasonality and trends) because of their fixed receptive fields. While some approaches reshape time series into 2D tensors based on frequency domain information (Wu et al., 2023) or downsample the time domain (Liu et al., 2022a), they fall short in comprehensively capturing long-range patterns. In contrast, transformer-based architectures (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2022b; Wang et al., 2022; Shi et al., 2024b) leverage token-wise self-attention to model long-range dependencies by allowing each token to attend to all others, overcoming the limitations of fixed receptive fields. However, unlike language tasks where tokens usually belong to distinct contexts, time series data often involve overlapping contexts at a single time point, such as daily, weekly, and seasonal patterns occurring simultaneously. This overlap makes it difficult to represent time series patterns effectively as tokens, posing challenges for transformer-based models in fully capturing the relevant temporal structures.
@@ 2 Related Work
Time Series Analysis. A pivotal aspect of time series analysis is the ability to extract diverse patterns from various time series while building powerful representations. This challenge has been explored across various model architectures. Traditional models like ARIMA (Anderson & Kendall, 1976) and STL (Cleveland et al., 1990) are effective for periodic and trend patterns but struggle with non-linear dynamics. Deep learning models, such as those by (Lai et al., 2018b) and (Zhao et al., 2017), capture sequential dependencies but face limitations with long-term dependencies. TCNs (Franceschi et al., 2019) improve local pattern extraction but are limited in capturing long-range dependencies. TimesNet (Wu et al., 2023) enhances long-range pattern extraction by treating time series as 2D signals, while MLP-based methods (Zeng et al., 2023; Ekambaram et al., 2023; Liu et al., 2023b; Wang et al., 2023c) offer simplicity and effectiveness. Transformer-based models like PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024) leverage self-attention to model long-range dependencies, demonstrating good forecasting performance. Given the strengths and limitations discussed above, there is a growing need for a TSPM capable of effectively extracting diverse patterns, adapting to various time series analytical tasks, and possessing strong generalization capabilities. As illustrated in Figure 1, TimeMixer++ meets this requirement by constructing robust representational capabilities, thereby demonstrating its potential for universal time series analysis.
@@
Hierarchical Time Series Modeling. Numerous methodologies have been advanced utilizing specialized deep learning architectures for time series analysis, with an emphasis on the decomposition and integration of temporal patterns. For example, several studies (Wu et al., 2021; Wang et al., 2024b; Zhou et al., 2022b; Luo et al., 2023; Wang et al., 2023b) utilize moving averages to discern seasonal and trend components, which are subsequently modeled using attention mechanisms (Wu et al., 2021; Zhou et al., 2022b; Shi et al., 2024a), convolutional networks (Wang et al., 2023a), or hierarchical MLP layers (Wang et al., 2024b; Wang, 2024). These components are individually processed prior to aggregation to yield the final output. Nonetheless, such approaches frequently depend on predefined and rigid operations for the disentanglement of seasonality and trends, thereby constraining their adaptability to complex and dynamic patterns. In contrast, as depicted in Figure 2, we propose a more flexible methodology that disentangles seasonality and trend directly within the latent space via dual-axis attention, thereby enhancing adaptability to a diverse range of time series patterns and task scenarios. Furthermore, by adopting a multi-scale, multi-resolution analytical framework (Mozer, 1991; Harti, 1993), we facilitate hierarchical interaction and integration across different scales and resolutions, substantially enhancing the effectiveness of time series modeling.
@@ Appendix A Implementation Details
To assess the effectiveness of our method across various tasks, we select 27 advanced baseline models spanning a wide range of architectures. Specifically, we utilize CNN-based models: MICN (2023a), SCINet (2022a), and TimesNet (2023); MLP-based models: TimeMixer (2024b), LightTS (2022a), and DLinear (2023); RMLP&RLinear (2023a) and Transformer-based models: iTransformer (2023a), PatchTST (2023), Crossformer (2023), FEDformer (2022b), Stationary (2022b), Autoformer (2021), and Informer (2021b). These models have demonstrated superior capabilities in temporal modeling and provide a robust framework for comparative analysis. For specific tasks, TiDE (2023a), FiLM (2022a), N-HiTS (2023), and N-BEATS (2019) address long- or short-term forecasting; Anomaly Transformer (2021) and MTS-Mixers (2023b) target anomaly detection; while Rocket (2023a), LSTNet (2018c), LSSL (2021), and Flowformer (2022) are utilized for classification. Few/zero-shot forecasting tasks employ ETSformer (2022), Reformer (2020), and LLMTime (2023).
@@ Appendix E Efficiency Analysis
We comprehensively compare the forecasting and imputation in performance, training speed, and memory footprint of the following models: TimeMixer++, iTransformer(Liu et al., 2023a), PatchTST(Nie et al., 2023), TimeMixer(Wang et al., 2024b), TIDE(Das et al., 2023a), Fedformer(Zhou et al., 2022b), TimesNet(Wu et al., 2023), MICN(Wang et al., 2023a), and SCINet(Liu et al., 2022a).
@@ Appendix G Error bars
Model TimeMixer++ iTransformer (2023a) Confidence

### v3

@@ 1 Introduction
At the core of TSPMs is their ability to recognize and generalize time series patterns inherent in time series data, enabling the model to uncover meaningful temporal structures and adapt to varying time series task scenarios. A line of research (Lai et al., 2018b; Zhao et al., 2017) has utilized recurrent neural networks (RNNs) to capture sequential patterns. However, these methods often struggle to capture long-term dependencies due to limitations like Markovian assumptions and inefficiencies. Temporal convolutional networks (TCNs) (Wu et al., 2023; Wang et al., 2023a; Liu et al., 2022a; Wang et al., 2024a) efficiently capture local patterns but face challenges with long-range dependencies (e.g., seasonality and trends) because of their fixed receptive fields. While some approaches reshape time series into 2D tensors based on frequency domain information (Wu et al., 2023) or downsample the time domain (Liu et al., 2022a), they fall short in comprehensively capturing long-range patterns. In contrast, transformer-based architectures (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2022b; Wang et al., 2022; Shi et al., 2024) leverage token-wise self-attention to model long-range dependencies by allowing each token to attend to all others, overcoming the limitations of fixed receptive fields. However, unlike language tasks where tokens usually belong to distinct contexts, time series data often involve overlapping contexts at a single time point, such as daily, weekly, and seasonal patterns occurring simultaneously. This overlap makes it difficult to represent time series patterns effectively as tokens, posing challenges for transformer-based models in fully capturing the relevant temporal structures.
@@ 2 Related Work
Time Series Analysis. A pivotal aspect of time series analysis is the ability to extract diverse patterns from various time series while building powerful representations. This challenge has been explored across various model architectures. Traditional models like ARIMA (Anderson & Kendall, 1976) and STL (Cleveland et al., 1990) are effective for periodic and trend patterns but struggle with non-linear dynamics. Deep learning models, such as those by (Lai et al., 2018b) and (Zhao et al., 2017), capture sequential dependencies but face limitations with long-term dependencies. TCNs (Franceschi et al., 2019) improve local pattern extraction but are limited in capturing long-range dependencies. TimesNet (Wu et al., 2023) enhances long-range pattern extraction by treating time series as 2D signals, while MLP-based methods (Zeng et al., 2023; Ekambaram et al., 2023; Liu et al., 2023; Wang et al., 2023c) offer simplicity and effectiveness. Transformer-based models like PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024) leverage self-attention to model long-range dependencies, demonstrating good forecasting performance. Given the strengths and limitations discussed above, there is a growing need for a TSPM capable of effectively extracting diverse patterns, adapting to various time series analytical tasks, and possessing strong generalization capabilities. As illustrated in Figure 1, TimeMixer++ meets this requirement by constructing robust representational capabilities, thereby demonstrating its potential for universal time series analysis.
@@
Hierarchical Time Series Modeling. Numerous methodologies have been advanced utilizing specialized deep learning architectures for time series analysis, with an emphasis on the decomposition and integration of temporal patterns. For example, several studies (Wu et al., 2021; Wang et al., 2024b; Zhou et al., 2022b; Luo et al., 2023; Wang et al., 2023b) utilize moving averages to discern seasonal and trend components, which are subsequently modeled using attention mechanisms (Wu et al., 2021; Zhou et al., 2022b; Shi et al., 2025), convolutional networks (Wang et al., 2023a), or hierarchical MLP layers (Wang et al., 2024b; Wang, 2024). These components are individually processed prior to aggregation to yield the final output. Nonetheless, such approaches frequently depend on predefined and rigid operations for the disentanglement of seasonality and trends, thereby constraining their adaptability to complex and dynamic patterns. In contrast, as depicted in Figure 2, we propose a more flexible methodology that disentangles seasonality and trend directly within the latent space via dual-axis attention, thereby enhancing adaptability to a diverse range of time series patterns and task scenarios. Furthermore, by adopting a multi-scale, multi-resolution analytical framework (Mozer, 1991; Harti, 1993), we facilitate hierarchical interaction and integration across different scales and resolutions, substantially enhancing the effectiveness of time series modeling.
@@ Appendix A Implementation Details
To assess the effectiveness of our method across various tasks, we select 27 advanced baseline models spanning a wide range of architectures. Specifically, we utilize CNN-based models: MICN (2023a), SCINet (2022a), and TimesNet (2023); MLP-based models: TimeMixer (2024b), LightTS (2022a), and DLinear (2023); RMLP&RLinear (2023a) and Transformer-based models: iTransformer (2024), PatchTST (2023), Crossformer (2023), FEDformer (2022b), Stationary (2022b), Autoformer (2021), and Informer (2021b). These models have demonstrated superior capabilities in temporal modeling and provide a robust framework for comparative analysis. For specific tasks, TiDE (2023a), FiLM (2022a), N-HiTS (2023), and N-BEATS (2019) address long- or short-term forecasting; Anomaly Transformer (2021) and MTS-Mixers (2023b) target anomaly detection; while Rocket (2023a), LSTNet (2018c), LSSL (2021), and Flowformer (2022) are utilized for classification. Few/zero-shot forecasting tasks employ ETSformer (2022), Reformer (2020), and LLMTime (2023).
@@ Appendix E Efficiency Analysis
We comprehensively compare the forecasting and imputation in performance, training speed, and memory footprint of the following models: TimeMixer++, iTransformer(Liu et al., 2024), PatchTST(Nie et al., 2023), TimeMixer(Wang et al., 2024b), TIDE(Das et al., 2023a), Fedformer(Zhou et al., 2022b), TimesNet(Wu et al., 2023), MICN(Wang et al., 2023a), and SCINet(Liu et al., 2022a).
@@ Appendix G Error bars
Model TimeMixer++ iTransformer (2024) Confidence